OpenStudy (anonymous):

Collect like terms 7x-9x=

OpenStudy (anonymous):

-2x

hero (hero):

I thought you knew how to do those already

OpenStudy (anonymous):

I barely

OpenStudy (anonymous):

\[7x-9x=x(7-9)=-2x\]

OpenStudy (anonymous):

it is when i see a variable I second guess it

hero (hero):

I'm sorry to hear that

OpenStudy (anonymous):

Whenever you see a variable, apply the distributive law.

OpenStudy (anonymous):

my distributive law knowledge sucks